### v1

# Distributed Scalable Coupled Policy Algorithm for Networked Multi-agent Reinforcement Learning

Abstract
This paper studies networked multi-agent reinforcement learning (NMARL) with interdependent rewards and coupled policies. In this setting, each agent’s reward depends on its own state-action pair as well as those of its direct neighbors, and each agent’s policy is parameterized by its local parameters together with those of its $\kappa_{p}$-hop neighbors, with $\kappa_{p}\geq 1$ denoting the coupled radius. The objective of the agents is to collaboratively optimize their policies to maximize the discounted average cumulative reward. To address the challenge of interdependent policies in collaborative optimization, we introduce a novel concept termed the “neighbors’ averaged $Q$-function” and derive a new expression for the coupled policy gradient. Based on these theoretical foundations, we develop a distributed scalable coupled policy (DSCP) algorithm, where each agent relies only on the state-action pairs of its $\kappa_{p}$-hop neighbors and the rewards its their $(\kappa_{p}+1)$-hop neighbors. Specially, in the DSCP algorithm, we employ a geometric 2-horizon sampling method that does not require storing a full $Q$-table to obtain an unbiased estimate of the coupled policy gradient. Moreover, each agent interacts exclusively with its direct neighbors to obtain accurate policy parameters, while maintaining local estimates of other agents’ parameters to execute its local policy and collect samples for optimization. These estimates and policy parameters are updated via a push-sum protocol, enabling distributed coordination of policy updates across the network. We prove that the joint policy produced by the proposed algorithm converges to a first-order stationary point of the objective function. Finally, the effectiveness of DSCP algorithm is demonstrated through simulations in a robot path planning environment, showing clear improvement over state-of-the-art methods.
Index Terms:
Networked multi-agent reinforcement learning, coupled policies, distributed scalable coupled policy algorithm.
I Introduction
With the advancement of artificial intelligence technology, reinforcement learning (RL) [1] has garnered increasing attention from both academia and industry. Currently, multi-agent reinforcement learning (MARL) has demonstrated outstanding performance on many complex scenarios, such as path planning [2, 3] and control and optimization system [4, 5, 6].
Network communication can substantially enhance multi-agent decision-making, which has led to the formulation of MARL over networks (i.e., networked MARL, NMARL). In the NMARL problem, agents leverage the network to exchange their local information (e.g., state-action, reward, and learned parameter) with their neighbors to achieve cooperative optimization.
I-A Related literature
Early studies primarily focused on fully connected interaction structures or assumed that agents had access to globally aggregated information. Notable examples include centralized training with decentralized execution frameworks, such as value-decomposition approaches (e.g., VDN [7], QMIX [8], QPLEX [9]), and cooperative actor-critic methods (e.g., COMA [10]). While these algorithms provide strong coordination performance, they rely on fully connected networks and centralized computation, which become increasingly impractical in large-scale systems.
To address scalability of the algorithm in large-scale systems, mean-field MARL approaches [11, 12] have recently been developed within the NMARL framework, leveraging communication networks. In these methods, each agent approximates the influence of others via the average behavior of its neighbors, thereby reducing dependence on global action information. Nevertheless, these approaches still require all agents to access a shared global state, which restricts their applicability in large-scale or communication-constrained environments.
To further reduce the algorithm’s dependence on the global state-action in the NMARL problem, a truncated $Q$-function technique was proposed in [13, 14], showing that each agent’s $Q$-function can be approximated using only the state-action pairs within its $\kappa$-hop neighborhood. Leveraging this property, several scalable algorithms were developed in [13, 14], where each agent relies solely on the state-action pairs of its $\kappa$-hop neighbors. Even so, the $Q$-functions are represented using a tabular format, which requires substantial storage as the communication range increases.
I-B Motivation
While the aforementioned algorithms have achieved significant progress in addressing the NMARL problem, they still face two major limitations. The first limitation arises from scalable NMARL methods that employ truncated $Q$-functions [13, 14]. In these approaches, the $Q$-functions are represented in tabular form, which introduces substantial storage requirements as the truncated range increases. Moreover, any approximation errors in the tabular $Q$-function will propagate to the policy gradient, potentially degrading the quality of the learned policy. The second limitation is that most existing algorithms generally assume that agents act according to independent policies. This assumption, however, is often unrealistic in practical scenarios, where agents’ decisions are inherently interdependent. In practical applications such as traffic signal control [15, 16], algorithms based on coupled policies typically exhibit significantly stronger learning capability than those relying on independent policies. The main reason is that the traffic dynamics at one intersection are tightly linked to the policies of neighboring intersections, and ignoring such coupling prevents intersections from collaboratively optimizing traffic flows. However, these algorithms for coupled policies have demonstrated promising empirical performance, but their theoretical foundations remain largely unexplored.
Motivated by these challenges, this paper addresses the following question: How to design a distributed and scalable algorithm for NMARL with coupled policies that avoids excessive computational overhead while ensuring convergence guarantees?
I-C Contribution
To address the aforementioned challenges, this paper makes the following key contributions.
(i) NMARL problem formulation with coupled policies: We formulate an NMARL framework with coupled policies, where each agent’s local policy is influenced not only by its own parameter but also by those of its $\kappa_{p}$-hop neighbors in a generalized manner. To eliminate the need for the global state-action information, a novel concept termed the “neighbors’ averaged $Q$-function” is introduced, based on which a new form of the coupled policy gradient is derived (see Theorem 2). Leveraging this gradient, we develop a distributed scalable coupled policy (DSCP) algorithm, where each agent relies only on the state-action pairs of its $\kappa_{p}$-hop neighbors and the rewards from its $(\kappa_{p}+1)$-hop neighbors.
(ii) Distributed algorithm with efficient sampling: In DSCP algorithm, we employ a geometric 2-horizon sampling method that does not require storing a full $Q$-table to obtain an unbiased estimate of the coupled policy gradient. Moreover, in the DSCP algorithm, each agent interacts only with its direct neighbors for true policy parameter and maintains its estimates of the policy parameters of other agents to execute its local policy and collect samples for optimization. These estimates and the policy parameters are then updated via a push-sum protocol, enabling distributed coordination of policy updates across the network.
(iii) Convergence guarantees: We establish that each agent optimizes its local policy parameters using an unbiased policy gradient estimate of the executed policy (see Lemma 2), and its local estimates of other agents’ policy parameters converge to their true values (see Theorem 3). Furthermore, the joint policy generated by the proposed DSCP algorithm converges to a first-order stationary point of the objective function (see Theorem 4).
The rest of this paper is organized as follows. The notations and preliminary concepts related to network are presented in Section II. The NMARL problem with coupled policies is described in Section III. Section IV presents the design of the DSCP algorithm. The convergence analysis of the proposed DSCP algorithm is provided in Section V. Simulation results demonstrating the effectiveness of the DSCP algorithm are presented in Section VI. Finally, the conclusion and further work are discussed in Section VII.
II Preliminaries
II-A Notations
$\mathbb{R}$ is the set of reals. $\mathbb{R}^{N}$ and $\mathbb{R}^{N\times M}$ represent the $N$-dimensional vector set and the $N\times M$-dimensional matrix set, respectively. For vectors $y\in\mathbb{R}^{N}, xx^{\top}$ denotes the transpose of $x$, and $x\otimes y$ denotes the Kronecker product of vectors $x$ and $y$. The operator $\|\cdot\|$ applied to vectors denotes the standard $\mathcal{L}_{2}$-norm, while applied to matrices denotes the induced $\mathcal{L}_{2}$-norm. The symbol $\mathbf{1}_{N}=(1,\cdots,1)^{\top}$ denotes the $N$-dimensional column vector consisting of all ones, whereas $\mathds{1}_{\{\cdot\}}$ represents the indicator function. $\mathbb{E}_{T_{1}}[\cdot]$ denotes the expectation taken over the random horizon length $T_{1}$. Similarly, $\mathbb{E}_{T_{1},T_{2}}[\cdot]$ represents the expectation with respect to the joint distribution of the random horizon lengths $T_{1}$ and $T_{2}$.
II-B Network
Denote $\mathcal{G}(\mathcal{N},\mathcal{E})$ as a directed network over $N$ agents, where $\mathcal{N}=\{1,\cdots,N\}$ is the set of agents and $\mathcal{E}$ is the set of edges among agents. An edge $e_{ij}=(j,i)\in\mathcal{E}$ means that agent $i$ can receive the information from agent $j$. For a given agent $i_{1}$ and agent $i_{k}$, if there exists an agent sequence $(i_{1},i_{2},\cdots,i_{k})$ such that $e_{i_{l+1},i_{l}}\in\mathcal{E}$ for all $1\leq l\leq k-1$, then the sequence $(i_{1},i_{2},\cdots,i_{k})$ is referred to as a directed path from agent $i_{1}$ to agent $i_{k}$. If for every edge $e_{i,j}\in\mathcal{E}$, there is always $e_{j,i}\in\mathcal{E}$, then the network is called an undirected network. An undirected network is considered connected if there exists at least one directed path from agent $i$ to agent $j$ for any $i,j\in\mathcal{N}$. Denote $\mathcal{N}_{i}=\{j|e_{ij}\in\mathcal{E},\forall j\in\mathcal{N}\}$ as the direct neighborhood of agent $i$ and $\mathcal{N}^{out}_{i}=\{j|e_{ji}\in\mathcal{E},\forall j\in\mathcal{N}\}$ as the out-neighborhood of agent $i$. For integer $\kappa_{p}\geq 1$, we let $\mathcal{N}^{\kappa_{p}}_{i}$ denote the $\kappa_{p}$-hop neighborhood of agent $i$, i.e., the agents whose graph distance to agent $i$ is less than or equal to $\kappa_{p}$, including $i$ itself. Moreover, we define $\mathcal{N}^{\kappa_{p}}_{-i}=\mathcal{N}\setminus\mathcal{N}^{\kappa_{p}}_{i}$ as a set of all agents except $\mathcal{N}^{\kappa_{p}}_{i}$. For any $i_{0}\in\mathcal{N}$, denote $\mathcal{N}^{\kappa_{p}}_{i,-i_{0}}=\mathcal{N}^{\kappa_{p}}_{i}\setminus\{i_{0}\}$ as the set of agent $i$’s $\kappa_{p}$-hop neighbors other than agent $i_{0}$. Denote $|\mathcal{N}^{out}_{i}|$ as the number of out-neighborhood of agent $i$ and $W=[w_{ij}]_{N\times N}$ as the weight matrix of $\mathcal{G}(\mathcal{N},\mathcal{E})$, which satisfies $w_{ij}=1/|\mathcal{N}^{out}_{j}|$ for $i\in\mathcal{N}^{out}_{j}$, otherwise $w_{ij}=0$.
III The NMARL problem with coupled policies
In this section, we present the formulation of the NMARL problem with coupled policies and introduce several preliminary results that will be used in the subsequent analysis.
III-A The NMARL problem setup
The model of the NMARL problem with coupled policies is described as $\big(\mathcal{G}(\mathcal{N},\mathcal{E}),\{\mathcal{S}_{i}\}_{i\in\mathcal{N}},\{\mathcal{A}_{i}\}_{i\in\mathcal{N}},$ $\{\mathcal{P}_{i}\}_{i\in\mathcal{N}},\{r_{i}\}_{i\in\mathcal{N}},\bm{\rho},\gamma\big)$, where $\mathcal{G}(\mathcal{N},\mathcal{E})$ is the underlying network among agents. $\mathcal{S}_{i}$ and $\mathcal{A}_{i}$ represent the local state space and the local action space of agent $i$, respectively. Let $s_{i}\in\mathcal{S}_{i}$ denote the local state and $a_{i}\in\mathcal{A}_{i}$ denote the local action of agent $i\in\mathcal{N}$. The global state $\bm{s}=(s_{1},\cdots,s_{N})$ and the global action $\bm{a}=(a_{1},\cdots,a_{N})$ are defined accordingly. Let $\bm{s}_{t}=(s_{1,t},\cdots,s_{N,t})$ denote the global state at time $t$, and let $\bm{a}_{t}=(a_{1,t},\cdots,a_{N,t})$ represent the global action at time $t$. Denote $s_{\mathcal{N}^{\kappa_{p}}_{i}}$ and $a_{\mathcal{N}^{\kappa_{p}}_{i}}$ as the states and actions of agent $i$’s $\kappa_{p}$-hop neighbors, respectively. Likewise, $s_{\mathcal{N}^{\kappa_{p}}_{i},t}$ and $a_{\mathcal{N}^{\kappa_{p}}_{i},t}$ represent the states and actions of agent $i$’s $\kappa_{p}$-hop neighbors at time $t$, respectively. $\mathcal{P}_{i}(s^{\prime}_{i}|s_{i},a_{i})$ is the state transition probability function of agent $i$. Define $\bm{\mathcal{P}}(\bm{s}^{\prime}|\bm{s},\bm{a})=\prod_{i=1}^{N}\mathcal{P}_{i}(s^{\prime}_{i}|s_{i},a_{i})$ as the global state transition probability function of agents. $r_{i}(s_{\mathcal{N}_{i}},a_{\mathcal{N}_{i}})$ is the reward function of agent $i$, which depends on the state-action pairs of its direct neighbors. For convenience, we define $r_{i,t}=r_{i}(s_{\mathcal{N}_{i},t},a_{\mathcal{N}_{i},t})$ as the instantaneous reward of agent $i$ at time $t$. $\bm{\rho}$ is the distribution of the initial state $\bm{s}_{0}$ and $\gamma\in(0,1)$ is the discount factor.
Remark 1
The NMARL formulation described above follows a standard framework that has been widely studied in the literature [17]. A representative motivating example is the distributed power control problem in wireless communication networks. In this setting, each communication link $i$ is modeled as an autonomous agent whose local state is typically given by its current transmission power level $0\leq p_{i}\leq p_{\mathrm{max}}$, where $p_{\mathrm{max}}$ denotes the maximum allowable power. The action $a_{i}\in\{0,-1,1\}$ indicates whether the agent maintains, decreases, or increases its transmission power by one unit. The local reward of agent $i$ is then defined as
$r_{i}(s_{\mathcal{N}_{i}},a_{\mathcal{N}_{i}})=\log\!\big(1+\frac{p_{i}g_{i,i}}{\sum_{j\in\mathcal{N}_{i}\setminus\{i\}}p_{j}g_{i,j}+\sigma_{i}}\big)-u_{i}p_{i},$ (1)
where $g_{i,j}$ denotes the channel gain from agent $j$ to agent $i$, $\sigma_{i}$ represents the noise power at agent $i$, and $u_{i}$ is a weighting parameter that balances the trade-off between throughput and power consumption. This model explicitly captures the interference coupling among neighboring communication links and illustrates the practical relevance of the proposed NMARL framework in wireless communication systems.
In light of the interplay among the decisions of agents, we formally define the local policy of agent $i$ as $\pi_{i}(\cdot|s_{i},\theta_{i},\theta_{\mathcal{N}^{\kappa_{p}}_{i,-i}})$, which depends not only on its local policy parameter but also incorporates the policy parameters of its $\kappa_{p}$-hop neighboring agents, where $\kappa_{p}\geq 1$ is constant. Denote $\bm{\pi}_{\bm{\theta}}(\cdot|\bm{s})\triangleq\bm{\pi}(\cdot|\bm{s},\bm{\theta})=\prod_{i=1}^{N}\pi_{i}(\cdot|s_{i},\theta_{i},\theta_{\mathcal{N}^{\kappa_{p}}_{i,-i}})$ as the joint policy of all agents, where $\bm{\theta}=(\theta^{\top}_{1},\cdots,\theta^{\top}_{N})^{\top}$ is the joint policy parameter of all agents.
In the NMARL problem, define the discounted average cumulative reward of agents under joint policy $\bm{\pi}_{\bm{\theta}}$ as
$J(\bm{\theta})=$ $\mathbb{E}_{\bm{s}\sim\bm{\rho}}\Big[\frac{1}{N}\!\sum_{t=0}^{\infty}\!\sum_{i=1}^{N}\gamma^{t}r_{i,t}|\bm{s}_{0}\!=\!\bm{s},\bm{a}_{t}\!\sim\!\bm{\pi_{\theta}}(\cdot|\bm{s}_{t})\Big].$ (2)
The goal of all agents is to find the optimal joint policy parameter to maximize $J(\bm{\theta})$, i.e., $\max_{\bm{\theta}}J(\bm{\theta})$.
For the purposes of the following development, we impose a structural assumption on the underlying network.
Assumption 1
In the NMARL problem, the underlying network $\mathcal{G}(\mathcal{N},\mathcal{E})$ is an undirected and connected network.
The assumption of undirected network is standard in NMARL [13, 14, 17] and is commonly adopted to characterize the communication relationships among agents.
III-B Preliminary results
In the NMARL problem, for any joint policy $\bm{\pi_{\theta}}$, define the global $Q$-function $Q^{\bm{\pi_{\theta}}}(\bm{s},\bm{a})$ as
$Q^{\bm{\pi_{\theta}}}(\bm{s},\bm{a})=\mathbb{E}_{\bm{\pi_{\theta}}}\Big[\frac{1}{N}\sum_{t=0}^{\infty}\sum_{i=1}^{N}\gamma^{t}r_{i,t}|\bm{s}_{0}=\bm{s},\bm{a}_{0}=\bm{a}\Big].$ (3)
and the local $Q$-function $Q^{\bm{\pi_{\theta}}}_{i}(s_{\mathcal{N}_{i}},a_{\mathcal{N}_{i}})$ of agent $i$ as
$Q^{\bm{\pi_{\theta}}}_{i}(s_{\mathcal{N}_{i}},a_{\mathcal{N}_{i}})=$ $\mathbb{E}_{\bm{\pi_{\theta}}}\Big[\sum_{t=0}^{\infty}\gamma^{t}r_{i,t}|s_{\mathcal{N}_{i},0}=s_{\mathcal{N}_{i}},a_{\mathcal{N}_{i},0}=a_{\mathcal{N}_{i}}\Big].$ (4)
Since the local reward function $r_{i}(s_{\mathcal{N}_{i}},a_{\mathcal{N}_{i}})$ of agent $i$ depends on the state-action pair $(s_{\mathcal{N}_{i}},a_{\mathcal{N}_{i}})$, the domain of its corresponding local $Q$-function in (4) naturally encompasses all possible combinations of $(s_{\mathcal{N}_{i}},a_{\mathcal{N}_{i}})$. Moreover, owing to the independence of the local state transition probabilities $\{\mathcal{P}_{i}(s^{\prime}_{i}|s_{i},a_{i})\}_{i\in\mathcal{N}}$, a decomposition of the global $Q$-function can be properly established in the following lemma.
Lemma 1
In the NMARL problem, for any joint policy $\bm{\pi_{\theta}}$, the global $Q$-function $Q^{\bm{\pi_{\theta}}}(\bm{s},\bm{a})$ can be decomposed as
$Q^{\bm{\pi_{\theta}}}(\bm{s},\bm{a})=\frac{1}{N}\sum_{i=1}^{N}Q^{\bm{\pi_{\theta}}}_{i}(s_{\mathcal{N}_{i}},a_{\mathcal{N}_{i}}).$ (5)
Lemma 1 achieves an exact decomposition of the global $Q$-function into the local $Q$-functions of individual agents. This decomposition is complete in the sense that it applies not only to the $Q$-values but also to the factorization of the global state-action space, which differs from the approach in [13, 14] and thus obviates the need for the truncation technique in the subsequent algorithm design.
Building upon the decomposition result in Lemma 1, we proceed to introduce the discounted state visitation distribution, a critical element in the derivation of the policy gradient theorem. Define $d^{\bm{\pi_{\theta}}}_{\bm{\rho}}(\bm{s})$ as the discounted state visitation distribution, given by
$d^{\bm{\pi_{\theta}}}_{\bm{\rho}}(\bm{s})=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\mathrm{Pr}^{\bm{\pi_{\theta}}}(\bm{s}_{t}=\bm{s}|\bm{s}_{0}\sim\bm{\rho}),$ (6)
where $\mathrm{Pr}^{\bm{\pi_{\theta}}}(\bm{s}_{t}=\bm{s}|\bm{s}_{0}\sim\bm{\rho})$ is the probability of occurrence of $\bm{s}_{t}=\bm{s}$ at time $t$ under the joint policy $\bm{\pi_{\theta}}$ and initial state distribution $\bm{\rho}$.
Based on $d^{\bm{\pi_{\theta}}}_{\bm{\rho}}(\bm{s})$ in (6), the coupled policy gradient $\nabla_{\theta_{i}}J(\bm{\theta})$ of agent $i$ is derived as follows.
Theorem 1
For any joint policy parameter $\bm{\theta}$ and agent $i\in\mathcal{N}$, the gradient of $J(\bm{\theta})$ with respect to $\theta_{i}$ is represented as
$\nabla_{\theta_{i}}J(\bm{\theta})=$ $\frac{1}{1-\gamma}\mathbb{E}_{\bm{s}\sim d^{\bm{\pi_{\theta}}}_{\bm{\rho}},\bm{a}\sim\bm{\pi_{\theta}}}\Big[\frac{1}{N}\sum_{l\in\mathcal{N}^{\kappa_{p}+1}_{i}}Q^{\bm{\pi_{\theta}}}_{l}(s_{\mathcal{N}_{l}},a_{\mathcal{N}_{l}})$
$\times\sum_{j\in\mathcal{N}^{\kappa_{p}}_{i}}\nabla_{\theta_{i}}\log\pi_{j}(a_{j}|s_{j},\theta_{j},\theta_{\mathcal{N}^{\kappa_{p}}_{j,-j}})\Big].$ (7)
Theorem 1 establishes the expression of the coupled policy gradient in the NMARL problem with coupled policies. It reveals that the coupled policy gradient $\nabla_{\theta_{i}}J(\bm{\theta})$ of agent $i$ in (1) depends on the local $Q$-functions $\{Q^{\bm{\pi_{\theta}}}_{l}(s_{\mathcal{N}_{l}},a_{\mathcal{N}_{l}})\}_{l\in\mathcal{N}^{\kappa_{p}+1}_{i}}$ from its $(\kappa_{p}+1)$-hop neighbors and the policies $\{\pi_{j}(a_{j}|s_{j},\theta_{j},\theta_{\mathcal{N}^{\kappa_{p}}_{j,-j}})\}_{j\in\mathcal{N}^{\kappa_{p}}_{i}}$ from its $\kappa_{p}$-hop neighbors.
By utilizing $\nabla_{\theta_{i}}J(\bm{\theta})$ in (1), the update of agent $i$’s local policy parameter $\theta_{i,t+1}$ can be formally expressed as
$\theta_{i,t+1}=\theta_{i,t}+\eta_{\theta,t}\nabla_{\theta_{i}}J(\bm{\theta}_{t}),$ (8)
where $\eta_{\theta,t}$ is the learning rate of policy parameter in $t$-th iteration of policy parameter.
Remark 2
It is important to observe that the iteration (8) requires each agent $i$ to utilize the exact local $Q$-functions associated with its $(\kappa_{p}+1)$-hop neighbors. For problems involving high-dimensional state-action spaces, the computation of these local $Q$-functions becomes computationally intractable, and the associated complexity grows exponentially with the number of agents. Although several existing works adopt tabular methods to estimate the $Q$-functions [13, 14], the approximation errors introduced by such estimation inevitably propagate to the policy update, thereby adversely affecting the learning performance. These limitations motivate the development of a distributed algorithm that avoids the explicit computation of the $Q$-functions or $Q$-tables, thereby enabling scalability in large-scale NMARL problems.
IV Distributed scalable coupled policy algorithm
To eliminate the need for agents to compute local $Q$-functions and improve the algorithm’s scalability, we first introduce a novel concept of the “neighbors’ averaged $Q$-function” and derive a new expression for coupled policy gradient. Based on this derivation, we propose a DSCP algorithm.
IV-A Neighbors’ averaged $Q$-function and a new expression of coupled policy gradient
From the definition of the local $Q$-function in (4), it can be inferred that the coupled policy gradient $\nabla_{\theta_{i}}J(\bm{\theta})$ of agent $i$ in (1) depends solely on the state-action pairs within its $(\kappa_{p}+2)$-hop neighbors and the rewards from its $(\kappa_{p}+1)$-hop neighbors. Motivated by this observation, we introduce a concept of “neighbors’ averaged $Q$-function” for each agent $i$, defined as
$\widehat{Q^{\bm{\pi_{\theta}}}_{i}}(s_{\mathcal{N}^{\kappa_{p}+2}_{i}},a_{\mathcal{N}^{\kappa_{p}+2}_{i}})$
$=$ $\mathbb{E}_{\bm{\pi_{\theta}}}\Big[\frac{1}{N}\sum^{\infty}_{t=0}\gamma^{t}\!\!\sum_{j\in\mathcal{N}^{\kappa_{p}+1}_{i}}\!\!r_{j}(s_{\mathcal{N}_{j},t},a_{\mathcal{N}_{j},t})\Big|$
$s_{\mathcal{N}^{\kappa_{p}+2}_{i},0}=s_{\mathcal{N}^{E,\kappa_{p}+2}_{i}},a_{\mathcal{N}^{\kappa_{p}+2}_{i},0}=a_{\mathcal{N}^{\kappa_{p}+2}_{i}}\Big].$ (9)
We emphasize that although the domain of $\widehat{Q^{\bm{\pi_{\theta}}}_{i}}(s_{\mathcal{N}^{\kappa_{p}+2}_{i}},a_{\mathcal{N}^{\kappa_{p}+2}_{i}})$ involves the state-action pairs of the $(\kappa_{p}+2)$-hop neighbors agent $i$, its value can be determined solely by the rewards from its $(\kappa_{p}+1)$-hop neighbors. This property arises from the specific structure of the local state transition function $\mathcal{P}_{i}(s^{\prime}_{i}|s_{i},a_{i})$, under which (IV-A) remains unaffected by the state-action pairs of agents outside the $(\kappa_{p}+2)$-hop neighborhood of agent $i$.
By using the neighbors’ averaged $Q$-function in (IV-A), a new expression for coupled policy gradient can be derived.
Theorem 2
For any joint policy $\bm{\pi_{\theta}}$, the gradient of $J(\bm{\theta})$ with respect to $\theta_{i}$ can be represented as
$\nabla_{\theta_{i}}J(\bm{\theta})=$ $\frac{1}{1-\gamma}\mathbb{E}_{\bm{s}\sim d^{\bm{\pi_{\theta}}}_{\bm{\rho}},\bm{a}\sim\bm{\pi_{\theta}}}\Big[\widehat{Q^{\bm{\pi_{\theta}}}_{i}}(s_{\mathcal{N}^{\kappa_{p}+2}_{i}},a_{\mathcal{N}^{\kappa_{p}+2}_{i}})$
$\times\sum_{j\in\mathcal{N}^{\kappa_{p}}_{i}}\nabla_{\theta_{i}}\log\pi_{j}(a_{j}|s_{j},\theta_{j},\theta_{\mathcal{N}^{\kappa_{p}}_{j,-j}})\Big].$ (10)
Theorem 2 fully exploits the notion of the neighbors’ averaged $Q$-function, providing an alternative formulation of the coupled policy gradient that explicitly incorporates the policies of agent $i$’s $\kappa_{p}$-hop neighbors, along with the term $\widehat{Q^{\bm{\pi_{\theta}}}_{i}}(s_{\mathcal{N}^{\kappa_{p}+2}_{i}},a_{\mathcal{N}^{\kappa_{p}+2}_{i}})$. Recalling the fact that each agent $i$ can compute the quantity of $\widehat{Q^{\bm{\pi_{\theta}}}_{i}}(s_{\mathcal{N}^{\kappa_{p}+2}_{i}},a_{\mathcal{N}^{\kappa_{p}+2}_{i}})$ in (IV-A) by using reward signals from its $(\kappa_{p}+1)$-hop neighbors. As a result, the policy gradient $\nabla_{\theta_{i}}J(\bm{\theta})$ in (2) for each agent $i$ can be computed solely from the state-action pairs and rewards within its limited neighborhood, thus providing a rigorous theoretical basis for the distributed algorithm design presented in the subsequent section.
IV-B DSCP algorithm
Before designing the DSCP algorithm, we introduce the following assumption concerning the collection of different types of information.
Assumption 2
In each iteration of policy parameter, every agent collects a single snapshot of the state-action pairs from its $\kappa_{p}$-hop neighbors and collects a trajectory of reward signals from its $(\kappa_{p}+1)$-hop neighbors. Moreover, the exchange of policy parameters is restricted to direct neighbors.
This assumption mirrors realistic communication constraints in networked multi-agent settings [19]: (i) State-action pairs tend to be high-volume and fast-changing, so a one-time snapshot per iteration is reasonable under bandwidth and latency limits; (ii) Reward signals, being lower in rate and lighter in data size, can be shared across a slightly larger neighborhood; (iii) Policy parameters frequently contain sensitive or strategically significant information. Restricting their exchange to only directly adjacent neighbors enhances confidentiality and reduces the risk of widespread dissemination.
In accordance with Theorem 2 and Assumption 2, the design of a distributed algorithm entails addressing two key issues: (i) Policy execution and (ii) Policy update.
For the issue of (i) Policy execution, under Assumption 2, each agent only receives the true policy parameters from its direct neighbors. When $\kappa_{p}\geq 2$, agents are unable to directly obtain the true policy parameters of all agents within their $\kappa_{p}$-hop neighborhood, which may hinder accurate policy execution. To address this limitation, we assume that each agent $i$ maintains an estimate of the policy parameters of other agents. Specifically, let $\hat{\theta}^{i}_{j,t}$ denote the local estimate of agent $i$ on the policy parameter of agent $j$ in $t$-th policy iteration. Inspired by the push-sum protocol in [20], the update for $\hat{\theta}^{i}_{j,t}$ is given by
$p_{i,t+1}=\sum_{l\in\mathcal{N}_{i}}w_{il}p_{l,t}$ (11a)
$\hat{\theta}^{i}_{j,t}\!=\!\frac{1}{p_{i,t+1}}\sum_{l\in\mathcal{N}_{i}}w_{il}\breve{\theta}^{l}_{j,t},$ (11b)
where $p_{i,1}=1$ for all $i\in\mathcal{N}$ and $\{\breve{\theta}^{i}_{j,t}\}_{i,j\in\mathcal{N}}$ denotes the intermediate variables, whose initial values satisfy $\breve{\theta}^{i}_{j,1}=\mathbf{0}_{d}$ for all $i,j\in\mathcal{N}$. Based on the estimated policy parameters $\{\hat{\theta}^{i}_{j,t}\}_{j\in\mathcal{N}}$, each agent $i$ executes its local policy as $\pi_{i}(a_{i}\mid s_{i},\hat{\theta}^{i}_{i,t},\hat{\theta}^{i}_{\mathcal{N}^{\kappa_{p}}_{i,-i},t})$ during the learning process.
For the issue of (ii) Policy update, it can be effectively addressed through the following three key steps.
Step 1: Define $\bm{\hat{\theta}}^{i}_{t}=\big((\hat{\theta}^{i}_{1,t})^{\top},\cdots,(\hat{\theta}^{i}_{N,t})^{\top}\big)^{\top}$, $\bm{\hat{\theta}}_{t}=\big((\bm{\hat{\theta}}^{1}_{t})^{\top},\cdots,(\bm{\hat{\theta}}^{N}_{t})^{\top}\big)^{\top}$, and $\bm{\hat{\pi}}_{\bm{\hat{\theta}}_{t}}=\prod_{i=1}^{N}\pi_{i}(a_{i}|s_{i},\hat{\theta}^{i}_{i,t},$ $\hat{\theta}^{i}_{\mathcal{N}^{\kappa_{p}}_{i,-i},t})$ as the executed joint policy in $t$-th iteration. Define $\nabla_{\theta_{i}}\tilde{J}(\bm{\hat{\theta}}_{t})$ as the coupled policy gradient under the joint policy $\bm{\hat{\pi}}_{\bm{\hat{\theta}}_{t}}$, which is represented as
$\nabla_{\theta_{i}}\tilde{J}(\bm{\hat{\theta}}_{t})=$ $\frac{1}{1-\gamma}\mathbb{E}_{\bm{s}\sim d^{\bm{\hat{\pi}}_{\bm{\hat{\theta}}_{t}}}_{\bm{\rho}},\bm{a}\sim\bm{\hat{\pi}}_{\bm{\hat{\theta}}_{t}}}\Big[\widehat{Q^{\bm{\hat{\pi}}_{\bm{\hat{\theta}}_{t}}}_{i}}(s_{\mathcal{N}^{\kappa_{p}+2}_{i}},a_{\mathcal{N}^{\kappa_{p}+2}_{i}})$
$\times\sum_{j\in\mathcal{N}^{\kappa_{p}}_{i}}\nabla_{\theta_{i}}\log\pi_{j}(a_{j}|s_{j},\hat{\theta}^{i}_{j,t},\hat{\theta}^{i}_{\mathcal{N}^{\kappa_{p}}_{j,-j},t})\Big]$ (12)
To estimate the policy gradient $\nabla_{\theta_{i}}\tilde{J}(\bm{\hat{\theta}}_{t})$ in (12), we employ a geometric 2-horizon sampling method. Specifically, we first select a random horizon $T_{1}\sim\mathrm{Geom}(1-\gamma)$ to generate a finite sample trajectory $(\bm{s},\bm{a})_{0:T_{1}}=(\bm{s}_{0},\bm{a}_{0},\bm{s}_{1},\cdots,\bm{s}_{T_{1}},\bm{a}_{T_{1}})$. According Assumption 2, each agent $i$ can collect $(s_{\mathcal{N}^{\kappa_{p}}_{i},T_{1}},a_{\mathcal{N}^{\kappa_{p}}_{i},T_{1}})$. Following this, we draw an additional horizon $T_{2}\sim\mathrm{Geom}(1-\gamma^{1/2})$ to generate another finite sample trajectory $(\bm{s},\bm{a})_{T_{1}:T_{1}+T_{2}}=(\bm{s}_{T_{1}},\bm{a}_{T_{1}},\bm{s}_{T_{1}+1},\cdots,\bm{s}_{T_{1}+T_{2}},\bm{a}_{T_{1}+T_{2}})$ for policy gradient estimation, where $T_{2}$ is statistically independent of $T_{1}$.
Step 2: By Assumption 2, each agent $i$ can collect a reward trajectory $\{r_{\mathcal{N}^{\kappa_{p}+1}_{i},t}\}_{t=T_{1}}^{T_{1}+T_{2}}$ that is associated with the sample trajectory $(\bm{s},\bm{a})_{T_{1}:T_{1}+T_{2}}$ to estimate the value of neighbors’ averaged $Q$-function $\widehat{Q^{\bm{\hat{\pi}}_{\bm{\hat{\theta}}_{t}}}_{i}}(s_{\mathcal{N}^{\kappa_{p}+2}_{i},T_{1}},a_{\mathcal{N}^{\kappa_{p}+2}_{i},T_{1}})$ in (IV-A) as
$\hat{Q}^{\bm{\hat{\pi}}_{\bm{\hat{\theta}}_{t}}}_{i,T_{1}}=\frac{1}{N}\sum_{\tau=0}^{T_{2}}\gamma^{\tau/2}\sum_{j\in\mathcal{N}^{\kappa_{p}+1}_{i}}r_{j,T_{1}+\tau}.$ (13)
Moreover, each agent $i$ can use the collected sample $(s_{\mathcal{N}^{\kappa_{p}}_{i},T_{1}},a_{\mathcal{N}^{\kappa_{p}}_{i},T_{1}})$ and its local policy parameter estimates $\bm{\hat{\theta}}^{i}_{t}$ to calculate
$\hat{g}^{\bm{\hat{\pi}}_{\bm{\hat{\theta}}_{t}}}_{i,T_{1}}\triangleq\sum_{j\in\mathcal{N}^{\kappa_{p}}_{i}}\nabla_{\theta_{i}}\log\pi_{j}(a_{j,T_{1}}|s_{j,T_{1}},\hat{\theta}^{i}_{j,t},\hat{\theta}^{i}_{\mathcal{N}^{\kappa_{p}}_{j,-j},t}).$ (14)
Based on (13) and (14), the coupled policy gradient estimation for $\nabla_{\theta_{i}}\tilde{J}(\bm{\hat{\theta}}_{t})$ in (12) can be calculated as
$\hat{\nabla}_{\theta_{i}}\tilde{J}(\bm{\hat{\theta}}_{t})=$ $\frac{1}{1-\gamma}\hat{Q}^{\bm{\hat{\pi}}_{\bm{\hat{\theta}}_{t}}}_{i,T_{1}}\hat{g}^{\bm{\hat{\pi}}_{\bm{\hat{\theta}}_{t}}}_{i,T_{1}}.$ (15)
Notably, $\hat{\nabla}_{\theta_{i}}\tilde{J}(\bm{\hat{\theta}}_{t})$ can also be regarded as an estimate of the true policy gradient $\nabla_{\theta_{i}}J(\bm{\theta}_{t})$ in (1), based on which the update of $\theta_{i,t+1}$ is formulated as
$\theta_{i,t+1}=\theta_{i,t}+\eta_{\theta,t}\hat{\nabla}_{\theta_{i}}\tilde{J}(\bm{\hat{\theta}}_{t}),$ (16)
where $\eta_{\theta,t}$ is the learning rate of policy parameter in $t$-th iteration.
Step 3: The update of the intermediate variables $\breve{\theta}^{i}_{j,t+1}$ of agent $i$ for $j$ is designed as
$\breve{\theta}^{i}_{j,t+1}=\left\{\begin{array}[]{ll}\sum\limits_{l\in\mathcal{N}_{i}}w_{il}\breve{\theta}^{l}_{j,t},~\mathrm{if}~j\notin\mathcal{N}_{i}\\ \sum\limits_{l\in\mathcal{N}_{i}}w_{il}\breve{\theta}^{l}_{j,t}+w_{ij}N(\theta_{j,t+1}-\theta_{j,t}),\mathrm{o/w}\end{array}\right$. (19)
where $w_{il}$ is the weight of the edge $e_{il}$. (19) guarantee the average property such that $(1/N)\sum_{i=1}^{N}\breve{\theta}^{i}_{j,t}=\theta_{j,t}$ for all $j\in\mathcal{N}$ and $t\geq 1$. By integrating (19), the updates in (11) ensure the average consensus convergence of $\hat{\theta}^{i}_{j,t}$ towards $\theta_{j,t}$.
Combining Steps 1-3, the detailed pseudo-code of the DSCP algorithm is presented in Algorithm 1.
1Input: The non-negative integers $T$, the learning-rate $\eta_{\theta,t}$, the initial policy parameters $\theta_{j,1}=\breve{\theta}^{i}_{j,1}=\mathbf{0}_{d}$ for all $i,j\in\mathcal{N}$;
2 for $t=1,2,\cdots,T-1$ do
3   Reset the initial state $\bm{s}_{0}\sim\bm{\rho}$ and draw $T_{1}\sim\mathrm{Geom}(1-\gamma)$ and $T_{2}\sim\mathrm{Geom}(1-\gamma^{1/2})$;
4   Each agent $i$ computes $\{\hat{\theta}^{i}_{j,t}\}_{j\in\mathcal{N}}$ by (11) and executes action $a_{i,0}\sim\pi_{i}(\cdot|s_{i,0},\hat{\theta}^{i}_{i,t},\hat{\theta}^{i}_{\mathcal{N}^{\kappa_{p}}_{i,-i},t})$;
5 for $\tau=1,2,\cdots,T_{1}$ do
6      Each agent $i$ obtains the local state $s_{i,\tau}$ and executes action $a_{i,\tau}\sim\pi_{i}(\cdot|s_{i,\tau},\hat{\theta}^{i}_{i,t},\hat{\theta}^{i}_{\mathcal{N}^{\kappa_{p}}_{i,-i},t})$;
7
8  Each agent $i$ collects $\{(s_{j,T_{1}},a_{j,T_{1}})\}_{j\in\mathcal{N}^{\kappa_{p}}_{i}}$ from its $\kappa_{p}$-hop neighbors and $\{r_{j,T_{1}}\}_{j\in\mathcal{N}^{\kappa_{p}+1}_{i}}$ from its $(\kappa_{p}+1)$-hop neighbors;
9 for $\tau=1,2,\cdots,T_{2}$ do
10      Each agent $i$ obtains the local state $s_{i,T_{1}+\tau}$ and executes action $a_{i,T_{1}+\tau}\!\!\sim\!\!\pi_{i}(\cdot|s_{i,T_{1}+\tau},\!\hat{\theta}^{i}_{i,t},\!\hat{\theta}^{i}_{\mathcal{N}^{\kappa_{p}}_{i,-i},t}\!)$;
11      Each agent $i$ collects $\{r_{j,T_{1}+\tau}\}_{j\in\mathcal{N}^{\kappa_{p}+1}_{i}}$ from its $(\kappa_{p}+1)$-hop neighbors;
12
13  Each agent $i$ calculates $\hat{\nabla}_{\theta_{i}}\tilde{J}(\bm{\hat{\theta}}_{t})$ by (15), and subsequently updates $\theta_{i,t+1}$ and $\{\breve{\theta}^{i}_{j,t+1}\}_{j\in\mathcal{N}}$ through (16) and (19), respectively;
14
Output: The joint policy $\bm{\pi}_{\bm{\theta}_{T}}$.
Algorithm 1 DSCP algorithm
Algorithm 1 entails the collection and exchange of two types of information among agents. (i) Environmental information collection: each agent $i$ collects state-action pairs $(s_{\mathcal{N}^{\kappa_{p}}_{i},T_{1}},a_{\mathcal{N}^{\kappa_{p}}_{i},T_{1}})$ for its $\kappa_{p}$-hop neighbors and rewards $\{r_{\mathcal{N}^{\kappa_{p}+1}_{i},t}\}_{t=T_{1}}^{T_{1}+T_{2}}$ from its $(\kappa_{p}+1)$-hop neighbors (see Lines 7 and 10). (ii) Parameter information exchange: each agent $i$ exchanges the true policy parameter $\theta_{i,t}$ and the intermediate variables $\{\breve{\theta}^{i}_{j,t}\}_{j\in\mathcal{N}}$ only with its direct neighbors (see Line 11).
Remark 3
In (11) and (19), each agent $i$ shares its true policy parameter $\theta_{i,t}$ only with its directly adjacent neighbors (see (19)). In addition, each agent $i$ exchanges the intermediate variables $\{\breve{\theta}^{i}_{j,t}\}_{j\in\mathcal{N}}$ with its direct neighbors (see (11b)). Note that these intermediate variables can be regarded as a form of obfuscated information, as they do not converge to the true parameter values in the same manner as the estimated policy parameters. This design prevents sensitive policy information from being directly propagated throughout the network, thereby enhancing confidentiality while still ensuring that the estimated parameters converge to their true values.
Remark 4
In contrast to $Q$-table-based methods [13, 14], which require each agent to maintain a complete $Q$-table encompassing all possible state-action pairs, the proposed Algorithm 1 obviates the necessity of storing a comprehensive $Q$-table. Instead, during each iteration, every agent $i$ computes only a single $Q$-value as defined in (13), which is subsequently utilized to construct the policy gradient estimate $\hat{\nabla}_{\theta_{i}}\tilde{J}(\bm{\hat{\theta}}_{t})$ in (15). Moreover, this estimate serves as an unbiased estimator of the policy gradient $\nabla_{\theta_{i}}\tilde{J}(\bm{\hat{\theta}}_{t})$ under the executed policy $\bm{\pi}_{\bm{\hat{\theta}}_{t}}$, as formally established in a subsequent Lemma 2.
Remark 5
It is worth noting that the proposed Algorithm 1 is designed for the case $\kappa_{p}\geq 2$. When $\kappa_{p}=1$, Assumption 2 implies that each agent can access the true policy parameters of its direct neighbors for executing policy $\bm{\pi}_{\bm{\theta}_{t}}$. In this scenario, the policy gradient $\nabla_{\theta_{i}}J(\bm{\theta}_{t})$ can be estimated by carrying out Steps 1 and 2 under the executed policy $\bm{\pi}_{\bm{\theta}_{t}}$, without relying on the additional parameter-estimation mechanism required for case where $\kappa_{p}\geq 2$.
V Convergence analysis
In this section, we begin by presenting several standard assumptions established in prior literature [13, 14, 21, 22], followed by the main theoretical results.
Assumption 3
For any agent $i\in\mathcal{N}$, there exists a constant $R>0$, such that the instantaneous reward $|r_{i}(s_{\mathcal{N}_{i}},a_{\mathcal{N}_{i}})|\leq R$ for all $(s_{\mathcal{N}_{i}},a_{\mathcal{N}_{i}})\in\mathcal{S}_{\mathcal{N}_{i}}\times\mathcal{A}_{\mathcal{N}_{i}}$.
Assumption 3 is standard in theoretical analyses of RL and indicates that the instantaneous rewards received by agents are bounded.
Assumption 4
For any joint policy $\bm{\pi_{\theta}}$, agent $i\in\mathcal{N}$, and agent $j\in\mathcal{N}^{\kappa_{p}}_{i}$, $\nabla_{\theta_{j}}\log\pi_{i}(a_{i}|s_{i},\theta_{i},\theta_{\mathcal{N}^{\kappa_{p}}_{i,-i}})$ exists and satisfies $\|\nabla_{\theta_{j}}\log\pi_{i}(a_{i}|s_{i},\theta_{i},\theta_{\mathcal{N}^{\kappa_{p}}_{i,-i}})\|\leq B$ with $B>0$ for any $(s_{i},a_{i})\in\mathcal{S}_{i}\times\mathcal{A}_{i}$. Moreover, for any agent $i\in\mathcal{N}$, $\nabla_{\theta_{i}}\log\pi_{i}(a_{i}|s_{i},\theta_{i},\theta_{\mathcal{N}^{\kappa_{p}}_{i,-i}})$ is $L$-Lipschitz continuous, i.e., for any $\bm{\theta},\bm{\theta}^{\prime}\in\mathbb{R}^{dN}$ and $(s_{i},a_{i})\in\mathcal{S}_{i}\times\mathcal{A}_{i}$, $\|\nabla_{\theta_{i}}\log\pi_{i}(a_{i}|s_{i},\theta_{i},\theta_{\mathcal{N}^{\kappa_{p}}_{i,-i}})-\nabla_{\theta_{i}}\log\pi_{i}(a_{i}|s_{i},\theta^{\prime}_{i},\theta^{\prime}_{\mathcal{N}^{\kappa_{p}}_{i,-i}})\|\leq L\|\bm{\theta}-\bm{\theta}^{\prime}\|$.
Assumption 4 implies that the policies of agents have bounded norms and exhibit Lipschitz continuity. In particular, this condition of policy function can be met by utilizing function classes such as softmax or log-linear functions.
Assumption 5
The learning rate $\eta_{\theta,t}$ in Algorithm 1 satisfies $\eta_{\theta,t}=\mathcal{O}(\frac{1}{t})$.
Assumption 5 is the standard assumption in the distributed optimization field and can ensure the convergence of estimated policy parameters to their true values.
V-A Intermediate lemmas
For notational convenience, for any $\kappa_{p}\geq 1$, we define
$M_{\kappa_{p}}=\max_{i\in\mathcal{N}}|\mathcal{N}^{\kappa_{p}}_{i}|,$ (20)
which characterizes the largest $\kappa_{p}$-hop neighborhood size among all agents. Recalling the definition of $\kappa_{p}$-hop neighborhoods in Section II-B, it follows that $1\leq M_{\kappa_{p}}\leq N$ for all $\kappa_{p}\geq 1$ and is non-decreasing with respect to $\kappa_{p}$.
Based on the above Assumptions 3-4, we have the following lemmas.
Lemma 2
Suppose Assumptions 3-4 hold. For the estimated policy parameter sequence $\{\bm{\hat{\theta}}_{t}\}_{1\leq t\leq T}$ generated by Algorithm 1, we have
(i) The estimated coupled policy gradient $\hat{\nabla}_{\theta_{i}}\tilde{J}(\bm{\hat{\theta}}_{t})$ is bounded, i.e., $\|\hat{\nabla}_{\theta_{i}}\tilde{J}(\bm{\hat{\theta}}_{t})\|\leq\hat{L}$ with $\hat{L}=\frac{BRM_{\kappa_{p}}M_{\kappa_{p}+1}}{(1-\gamma)(1-\gamma^{1/2})N}$;
(ii) $\hat{\nabla}_{\theta_{i}}\tilde{J}(\bm{\hat{\theta}}_{t})$ is an unbiased estimate of $\nabla_{\theta_{i}}\tilde{J}(\bm{\hat{\theta}}_{t})$, i.e., $\mathbb{E}_{T_{1},T_{2}}[\hat{\nabla}_{\theta_{i}}\tilde{J}(\bm{\hat{\theta}}_{t})|\bm{\hat{\theta}}_{t}]=\nabla_{\theta_{i}}\tilde{J}(\bm{\hat{\theta}}_{t})$.
The first part establishes that the norm of the estimated policy gradient $\hat{\nabla}_{\theta_{i}}\tilde{J}(\bm{\hat{\theta}}_{t})$ is upper-bounded, with the bound explicitly dependent on the choice of $\kappa_{p}$. Moreover, the second part demonstrates that Algorithm 1 provides each agent $i$ with an unbiased estimate of the coupled policy gradient under the executed joint policy $\bm{\hat{\pi}}_{\bm{\hat{\theta}}_{t}}$ during the $t$-th iteration, which is essential for establishing the convergence of Algorithm 1 in the subsequent Theorem 4.
Let $L_{1}=\big(\frac{LRM_{\kappa_{p}}M_{\kappa_{p}+1}}{(1-\gamma)^{2}N}+\frac{(1+\gamma)B^{2}RM^{2}_{\kappa_{p}}M_{\kappa_{p}+1}}{(1-\gamma)^{3}N^{\frac{1}{2}}}\big)$ and $L_{2}=\big(\frac{LRM_{\kappa_{p}}M_{\kappa_{p}+1}}{(1-\gamma)^{2}N^{\frac{1}{2}}}+\frac{(1+\gamma)B^{2}RM^{2}_{\kappa_{p}}M_{\kappa_{p}+1}}{(1-\gamma)^{3}}\big)$, we have the following results.
Lemma 3
Suppose Assumptions 3-4 hold. For any agent $i\in\mathcal{N}$ and joint policy parameters $\bm{\theta},\bm{\theta}^{\prime}\in\mathbb{R}^{dN}$, we have
$\|\nabla_{\theta_{i}}J(\bm{\theta})-\nabla_{\theta_{i}}J(\bm{\theta}^{\prime})\|\leq L_{1}\|\bm{\theta}-\bm{\theta}^{\prime}\|,$ (21)
which also means that $J(\bm{\theta})$ is $L_{2}$-smooth.
Lemma 3 proves that under the Assumptions 3-4, the objective function $J(\bm{\theta})$ with coupled policy in (2) satisfies the $L_{2}$-smooth, with the constant $L_{2}$ depending explicitly on $\kappa_{p}$.
[…]
Corollary 1
Suppose Assumptions 3-4 hold. For any agent $i\in\mathcal{N}$, any joint policies $\bm{\pi_{\theta}}$ and $\bm{\pi}_{\bm{\hat{\theta}}}$, we have $\nabla_{\theta_{i}}J(\bm{\theta})=\nabla_{\theta_{i}}\tilde{J}(\mathbf{1}_{N}\otimes\bm{\theta})$ and
$\|\nabla_{\theta_{i}}\tilde{J}(\mathbf{1}_{N}\otimes\bm{\theta})-\nabla_{\theta_{i}}\tilde{J}(\bm{\hat{\theta}})\|\leq L_{1}\|\mathbf{1}_{N}\otimes\bm{\theta}-\bm{\hat{\theta}}\|$. (22)
Corollary 1 closely parallels that of Lemma 3. It characterizes the error between the policy gradient under executed joint policy $\bm{\hat{\pi}}_{\bm{\hat{\theta}}}$ and that under the true joint policy $\bm{\pi_{\theta}}$.
V-B Main results
The main results, such as the convergence of the policy parameter estimation and first-order stationary convergence of Algorithm 1, are presented as below.
Theorem 3
Suppose Assumptions 1-5 hold. In Algorithm 1, for all $i,j\in\mathcal{N}$, the local policy parameter estimate $\hat{\theta}^{i}_{j,t}$ of agent $i$ with respect to agent $j$ converges to agent $j$’s true policy parameter $\theta_{j,t}$ with the rate $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{O}(\frac{M_{\kappa_{p}}M_{\kappa_{p}+1}\log{t}}{t})$}, i.e.,
$\|\hat{\theta}^{i}_{j,t}-\theta_{j,t}\|={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{O}(\frac{M_{\kappa_{p}}M_{\kappa_{p}+1}\log{t}}{t})},\forall i,j\in\mathcal{N}.$ (23)
Theorem 3 establishes that the policy parameter estimation $\hat{\theta}^{i}_{j,t}$ of agent $i$ for agent $j$ converges to the true policy parameter $\theta_{j,t}$ as $t\rightarrow\infty$. Moreover, the convergence rate depends on the choice of the coupled radius $\kappa_{p}$. As the number of iterations increases, Theorem 3 ensures that the agents can generate reliable sample data based on the executed joint policy $\bm{\hat{\pi}}_{\bm{\hat{\theta}}_{t}}$, thereby facilitating the cooperative optimization of policy parameters.
Theorem 4
Suppose Assumptions 1-5 hold. There exists a positive constant $J_{\mathrm{sup}}(M_{\kappa_{p}},M_{\kappa_{p}+1})>0$, which depends on the neighborhood sizes $M_{\kappa_{p}}$ and $M_{\kappa_{p}+1}$ and is monotonically increasing in both arguments, such that the policy parameter sequence $\{\bm{\theta}_{t}\}_{t=1}^{T-1}$ generated by Algorithm 1 satisfies $\frac{\mathbb{E}\big[\sum_{t=1}^{T-1}\eta_{\theta,t}\|\nabla_{\bm{\theta}}J(\bm{\theta}_{t})\|^{2}\big]}{\sum_{t=1}^{T-1}\eta_{\theta,t}}\leq\frac{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}J_{\mathrm{sup}}(M_{\kappa_{p}},M_{\kappa_{p}+1})}}{\sum_{t=1}^{T-1}\eta_{\theta,t}}$ and
$\lim_{T\rightarrow\infty}\frac{\mathbb{E}\Big[\sum_{t=1}^{T-1}\eta_{\theta,t}\|\nabla_{\bm{\theta}}J(\bm{\theta}_{t})\|^{2}\Big]}{\sum_{t=1}^{T-1}\eta_{\theta,t}}=0.$ (24)
Theorem 4 indicates that Algorithm 1 converges to a first-order stationary point of the objective function. In addition, the convergence rate is inversely related to the value of $\kappa_{p}$, and a larger $\kappa_{p}$ typically results in slower convergence.
V-C Extend results for reward function with larger coupled radius
Our proposed framework can also extend to settings where each agent’s reward depends on a larger neighborhood, under the following assumption.
Assumption 6
The reward function of agent $i$ adopts the form $r_{i}(s_{\mathcal{N}^{\kappa_{r}}_{i}},a_{\mathcal{N}^{\kappa_{r}}_{i}})$, where $1\leq\kappa_{r}\leq\kappa_{p}$ represents a coupled radius that depicts the reward dependency range. For any agent $i\in\mathcal{N}$, there exists a constant $R>0$, such that the instantaneous reward $|r_{i}(s_{\mathcal{N}^{\kappa_{r}}_{i}},a_{\mathcal{N}^{\kappa_{r}}_{i}})|\leq R$ for all $(s_{\mathcal{N}^{\kappa_{r}}_{i}},a_{\mathcal{N}^{\kappa_{r}}_{i}})\in\mathcal{S}_{\mathcal{N}^{\kappa_{r}}_{i}}\times\mathcal{A}_{\mathcal{N}^{\kappa_{r}}_{i}}$.
Under Assumption 6, the main results in Section V-B remain valid after adjusting the locality radii that appear in several key expressions. Specifically, the radius $(\kappa_{p}+1)$ in (1), (IV-A), and (13) is replaced by $(\kappa_{p}+\kappa_{r})$, and the radius $(\kappa_{p}+2)$ in (IV-A), (2), and (12) is replaced by $(\kappa_{p}+2\kappa_{r})$. Furthermore, Algorithm 1 requires redevelopment under the following assumption, which is a slightly modified version of Assumption 2.
Assumption 7
In each iteration of policy parameter, every agent collects a single snapshot of the state-action pairs from its $\kappa_{p}$-hop neighbors and collects a trajectory of reward signals from its $(\kappa_{p}+\kappa_{r})$-hop neighbors. Moreover, the exchange of policy parameters is restricted to direct neighbors.
Based on Assumption 7, Algorithm 1 can be extended in a direct manner, yielding the following results.
Corollary 2
Suppose Assumptions 1 and 4-7 hold. In the extend Algorithm 1 for $\{r_{i}(s_{\mathcal{N}^{\kappa_{r}}_{i}},a_{\mathcal{N}^{\kappa_{r}}_{i}})\}_{i\in\mathcal{N}}$, the local policy parameter estimate $\hat{\theta}^{i}_{j,t}$ of agent $i$ with respect to agent $j$ converges to agent $j$’s true policy parameter $\theta_{j,t}$ with a rate $\mathcal{O}(\frac{M_{\kappa_{p}}M_{\kappa_{p}+\kappa_{r}}\log{t}}{t})$. Moreover, there exists a positive constant $J_{\mathrm{sup}}(M_{\kappa_{p}},M_{\kappa_{p}+\kappa_{r}})>0$ that is employed to substitute $J_{\mathrm{sup}}(M_{\kappa_{p}},M_{\kappa_{p}+1})$ in Theorem 4 to guarantee its validity.
Corollary 2 directly extends Theorems 3 and 4, demonstrating our DSCP algorithm’s generalizability to different coupled ranges in the reward function.
VI Simulation
This section evaluates Algorithm 1 in a robot swarm path-planning environment. Although a related setup appears in [2], the task considered here is significantly more complex.
VI-A The description of the path planning problem
The path planning problem of $N=10$ robots (i.e., agents) on the acyclic path structure (as shown in Fig. 1) is considered, where the “blue” nodes $\{b_{1},b_{2},\cdots,b_{5}\}$ represent the set of starting locations for agents and the “red” node represents the destination location.
Figure 1: Path structure and underlying network.
We formulate a NMARL model $\big({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{G}(\mathcal{N},\mathcal{E})},\{\mathcal{S}_{i}\}_{i\in\mathcal{N}},$ $\{\mathcal{A}_{i}\}_{i\in\mathcal{N}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\{\mathcal{P}^{s}_{i}\}_{i\in\mathcal{N}}},\{r_{i}\}_{i\in\mathcal{N}},\gamma\big)$ for the path planning problem, where $\mathcal{G}$ denotes the underlying network among agents with $\mathcal{N}=\{1,\cdots,N\}$ representing the set of all agents. The detailed structure of the underlying network is elaborated in Fig. 1. $\mathcal{S}_{i}=\{b_{1},b_{2},b_{3},b_{4},b_{5},c_{1},c_{2},c_{3},c_{4},d_{1},d_{2},d_{3},e\}$ denotes the local state space of agent $i\in\mathcal{N}$, and $\mathcal{A}_{i}=\{0,1,2\}$ denotes its local action space. In this problem, the state transition $s^{\prime}_{i}=\mathcal{P}^{s}_{i}(s_{i},a_{i})$ of agent $i$ is governed by the following rules: (i) The local action “0” means that agent $i$ will remain stationary at the current location for one step; (ii) If the out-degree of the current location is 2, the actions “1” and “2” indicate that agent $i$ will follow the upper edge and the lower edge, respectively; (iii) If the action selected by agent $i$ exceeds the out degree of the current location, it will remain at the current location for one step. Let $P^{s_{i},a_{i}}_{i}=\big(s_{i}\rightarrow\mathcal{P}^{s}_{i}(s_{i},a_{i})\big)$ denote the movement of agent $i$, the reward function $r_{i}(s_{\mathcal{N}_{i}},a_{\mathcal{N}_{i}})$ of agent $i$ is defined as
$r_{i}(s_{\mathcal{N}_{i}},a_{\mathcal{N}_{i}})=\left\{\begin{array}[]{ll}-r_{\epsilon},~\mathrm{if}~s_{i}=\mathcal{P}^{s}_{i}(s_{i},a_{i})\\ -r_{\epsilon}-\frac{0.5*|\{j\in\mathcal{N}_{i}|P^{s_{i},a_{i}}_{i}=P^{s_{j},a_{j}}_{j}\}|}{N},\mathrm{o/w},\end{array}\right$. (27)
where the first term $r_{\epsilon}=0.5$ denotes the time cost incurred at each step, while the second term captures the collision penalty when agent $i$ shares the same path with its neighbors. Moreover, $|\{j\in\mathcal{N}_{i}\mid P^{s_{i},a_{i}}_{i}=P^{s_{j},a_{j}}_{j}\}|$ represents the number of neighbor agents that share the same path with agent $i$.
In this path planning problem, we set the initial locations of agents as $\{b_{1},b_{2},b_{3},b_{4},b_{5},b_{1},b_{2},b_{3},b_{4},b_{5}\}$ and the discount factor as $\gamma=0.9$. The objective of agents is to reach the destination as quickly as possible while minimizing collisions, i.e., maximize the objective function in (2).
VI-B The results of Algorithm 1
In Algorithm 1, the local coupled policy $\pi_{i}(a_{i}|s_{i},\theta_{i},\theta_{\mathcal{N}^{\kappa_{p}}_{i,-i}})$ for agent $i$ is defined as
$\pi_{i}(a_{i}|s_{i},\theta_{i},\theta_{\mathcal{N}^{\kappa_{p}}_{i,-i}})$
$=$ $\frac{\exp\big(0.9*\theta_{i,s_{i},a_{i}}+(0.1/|\mathcal{N}^{\kappa_{p}}_{i,-i}|)\sum_{j\in\mathcal{N}^{\kappa_{p}}_{i,-i}}\theta_{j,s_{i},a_{i}}\big)}{\sum_{a^{\prime}_{i}}\exp\big(0.9*\theta_{i,s_{i},a^{\prime}_{i}}+(0.1/|\mathcal{N}^{\kappa_{p}}_{i,-i}|)\sum_{j\in\mathcal{N}^{\kappa_{p}}_{i,-i}}\theta_{j,s_{i},a^{\prime}_{i}}\big)},$
where $\theta_{i}\in\mathbb{R}^{|\mathcal{S}_{i}||\mathcal{A}_{i}|}$ for all $i\in\mathcal{N}$.
VI-B1 Performance of Algorithm 1
To demonstrate the advantages of our Algorithm 1, we compared it with the state-of-the-art scalable actor-critic (SAC) algorithm [13] and distributed policy gradient (DPG) algorithm [17] under multiple random seeds.
Figure 2: The evolution of $J(\bm{\theta}_{t})$ produced by Algorithm 1, SAC algorithm, and DPG algorithm.
The evolution of objective function $J(\bm{\theta}_{t})$ produced by Algorithm 1 with $\kappa_{p}=2$, SAC algorithm, and DPG algorithm are shown in Fig. 2, where the solid line represents the average value of $J(\bm{\theta}_{t})$ over multiple random seeds and the shaded area represents its variance. As shown in Fig. 2, Algorithm 1 consistently outperforms both the SAC and DPG algorithms, exhibiting a notably faster convergence rate. These results demonstrate that the coupled policy framework provides enhanced learning capabilities compared to the independent policy approaches in SAC algorithm and DPG algorithm. We also observe that the DPG algorithm converges more slowly, which can be attributed to the fact that it directly estimates the state-value function rather than the $Q$-function.
(a) $\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\|\hat{\theta}^{i}_{j,t}-\theta_{j,t}\|$
(b) $\|\nabla_{\bm{\theta}}J(\bm{\theta}_{t})\|$
Figure 3: The evolution of the policy parameter estimation error and the norm of policy gradient produced by Algorithm 1.
Fig. 3 shows the evolution of the policy parameter estimation error $(1/N^{2})\sum_{i=1}^{N}\sum_{j=1}^{N}\|\hat{\theta}^{i}_{j,t}-\theta_{j,t}\|$ and the norm of policy gradient $\|\nabla_{\bm{\theta}}J(\bm{\theta}_{t})\|$ produced by Algorithm 1. As illustrated in Fig. 3(a), the policy parameter estimation error converges to 0, confirming the convergence of $\hat{\theta}^{i}_{j,t}$ to $\theta_{j,t}$ as stated in Theorem 3. Furthermore, Fig. 3(b) shows that the norm of the policy gradient, $\|\nabla_{\bm{\theta}}J(\bm{\theta}_{t})\|$, also converges to zero during the learning process, indicating that Algorithm 1 converges to a stationary point of $J(\bm{\theta})$ (i.e., Theorem 4).
VI-B2 Ablation experiment
An ablation study is conducted by setting $\kappa_{p}=0,1,2$ in Algorithm 1. Fig. 4 illustrates the influence of the coupled radius $\kappa_{p}$ on both the convergence rate and the steady-state performance of Algorithm 1.
Figure 4: The performances of Algorithm 1 with different $\kappa_{p}$ on objective function.
Convergence rate: In Fig. 4, when $\kappa_{p}=0$, the algorithm exhibits the fastest convergence, reaching a stable point at approximately $5.6\times 10^{3}$ iterations. However, the faster convergence is accompanied by noticeably degraded solution quality, as the policy with $\kappa_{p}=0$ is learned in a fully independent manner and does not exploit any neighborhood information. For $\kappa_{p}=1$ and $\kappa_{p}=2$, the algorithm converges at approximately $1.2\times 10^{4}$ and $2.0\times 10^{4}$ iterations, respectively. This behavior aligns with the result in Theorem 4, which states that the convergence rate of Algorithm 1 decreases with increasing $\kappa_{p}$, owing to expanded local policy dependencies and the associated increase in estimation complexity.
Steady-state performance: As shown in Fig. 4, the steady-state performance of Algorithm 1 improves monotonically with increasing $\kappa_{p}$. Notably, the performance improvement from $\kappa_{p}=0$ to $\kappa_{p}=1$ is substantial, whereas the gain from $\kappa_{p}=1$ to $\kappa_{p}=2$ is relatively marginal. This is because, under the given simulation setup, one-hop coupling can already capture the primary aspects of coordination among agents.
VII Conclusions
In this paper, we proposed a novel DSCP algorithm for NMARL problem with coupled policies and proved that it reaches a stationary point of the objective function. The foundational framework of the DSCP algorithm incorporates the neighbors’ averaged $Q$-function and the distributed optimization of policy parameters. In future work, the use of function approximation techniques for neighbors’ averaged $Q$-function in coupled-policy settings will be explored to further improve learning efficiency.